OpenStudy (anonymous):

John has a jar filled with 150 coins made up of quarters and nickels. The total value of the coins is $26.70. How many quarters and how many nickels are in the jar?

OpenStudy (anonymous):

answers could be 75 quarters; 75 nickels 96 quarters; 54 nickels 104 quarters; 46 nickels 54 quarters; 96 nickels

OpenStudy (anonymous):

@Dobby1

OpenStudy (anonymous):

@Dobby1 @Easyaspi314 help!!

OpenStudy (dobby1):

hey sorry I am late

OpenStudy (anonymous):

its cool i already solved this one here is another

OpenStudy (anonymous):

Faith mixes 14 liters of 12% acid solution with a 30% acid solution which results in an 18% acid solution. If x represents the number of liters of 30% acid solution used, which equation can be used to find the number of liters of 30% acid solution in the mixture?

OpenStudy (anonymous):

answers could be 1.68 + 0.3x = 0.18(x + 14) 2.52 + 0.3x = 0.12(x + 14) 4.2 + 0.18x = 0.3(x + 14) 1.68 + 0.18x = 0.3(x + 14)

OpenStudy (dobby1):

I would honestly pick the top one because the answer should have to do with 30% or .3 times the number of liters = 18% or .18 times the liters of 30% acid solution and 12%solution so I think A is the answer. :)

OpenStudy (anonymous):

Marie mixes 16 liters of 18% acid solution with a 27% acid solution which results in 21% acid solution. here is another one Find the number of liters of 27% solution in the new mixture

OpenStudy (anonymous):

answers could be 8 liters 12 liters 13.7 liters 16 liter

OpenStudy (dobby1):

do they tell you the number of 21 percent acid solution their is

OpenStudy (anonymous):

i dont think so no

OpenStudy (dobby1):

ok looks like we are using the equation we found above this is going to be fun

OpenStudy (anonymous):

ok lets make it fun!

OpenStudy (dobby1):

Ok well let me post how the equation is going to look

OpenStudy (dobby1):

\[.18(16)+.27x=.21(x+16)\]

OpenStudy (dobby1):

first we have to multiply the stuff in the parenthasese by what is right next to is

OpenStudy (dobby1):

so that means what we are multiplying is .18*16 and then we have .21*x+.21*16

OpenStudy (dobby1):

so what is .18*16

OpenStudy (dobby1):

you can put it in a calculator

OpenStudy (dobby1):

@manguitar you still there?

OpenStudy (anonymous):

yes im sorry

OpenStudy (dobby1):

its fine just tell me what .18*16 equals

OpenStudy (anonymous):

1.2 @Dobby1 ?

OpenStudy (anonymous):

2.88

OpenStudy (dobby1):

2.88 actually and brb afk

OpenStudy (anonymous):

so what would that be in liters

OpenStudy (dobby1):

no the 16 was liters we are supposed to multipy them together

OpenStudy (anonymous):

so what would the answer be? out of the ones i gave you?

OpenStudy (dobby1):

we are not done quite yet to find the answer we have to isolate x

OpenStudy (anonymous):

ok

OpenStudy (dobby1):

so now our equation looks like \[2.88+27x=.21(x+16)\]

OpenStudy (anonymous):

ok continue

OpenStudy (dobby1):

the next step is to multiply the x on the right side of the equal sign by .21 giving us .21x

OpenStudy (anonymous):

next

OpenStudy (dobby1):

that gives us \[2.88+27x=.21x+(.21*16)\]

OpenStudy (anonymous):

ok next

OpenStudy (dobby1):

now what I want you to do for this step is use a calculator to multiply .21*16

OpenStudy (anonymous):

3.36

OpenStudy (dobby1):

good now our equation looks like \[2.88+27x=.21x+3.36\]

OpenStudy (anonymous):

ok next

OpenStudy (anonymous):

= 3.57z

OpenStudy (anonymous):

@gorv we are working on a different problem

OpenStudy (dobby1):

so now we have to get rid of the .21x my subtracting it from both sides so now use a calculator and do this problem 27-.21

OpenStudy (anonymous):

26.79

OpenStudy (dobby1):

good @manguitar

OpenStudy (anonymous):

its all good thank for helping though @gorv

OpenStudy (dobby1):

\[2.88+26.79x=3.36\]

OpenStudy (dobby1):

now I want you to do the following equation in your calculator 3.36-2.88

OpenStudy (anonymous):

0.48

OpenStudy (dobby1):

good now our equation is \[26.79x=.48\]

OpenStudy (anonymous):

ok next

OpenStudy (dobby1):

ok now you need to put the following equation into your calculator .48/26.79

OpenStudy (anonymous):

0.01791713325

OpenStudy (dobby1):

hmm

OpenStudy (dobby1):

ok new plan

OpenStudy (dobby1):

we are going to plug in the answer choices they gave you that should work

OpenStudy (anonymous):

ok good

OpenStudy (dobby1):

because the answer we got just now is way wrong so here are the answers

OpenStudy (dobby1):

\[.18(16)+.27x=.21(x+16)\] 8 liters 12 liters 13.7 liters 16 liter

OpenStudy (dobby1):

to solve we are going to have to plug each answer in for x

OpenStudy (anonymous):

ok

OpenStudy (dobby1):

so for the first answer our equation looks like \[.18(16)+.27(8)=.21(8+16)\]

OpenStudy (anonymous):

ok next

OpenStudy (dobby1):

when we do all the multiplying we get \[2.88+2.16=1.68+3.36\]

OpenStudy (dobby1):

then we have to add and that gives us \[5.04=5.04\]

OpenStudy (anonymous):

that seems right

OpenStudy (dobby1):

ok it looks like we are right the answer is 8 liters I would love to help you more but I have to get some school done and eat lunch

OpenStudy (anonymous):

ok me too thanks a lot! @Dobby1